OpenStudy (anonymous):

Add: (4s^4-3s^3+5s^2+17s-8)+(s^5+9s^3+6s^2-5s+3) + (-7s^4+s^2-7s-4)

OpenStudy (anonymous):

ok.. lets add all the like terms together then string 'em all together in the end... the highest degree is 5 and you only have one of 'em . next is s^4: 4s^4 + -7s^4 = s^3: -3s^3 + 9s^3 = s^2: 5s^2 + 6s^2 + s^2 = s^1: 17s + -5s + -7s = s^0: -8 + 3 + -4 = can you find all these?

OpenStudy (anonymous):

s^4:-3s^4 s^3:6s^3 s^2: 12s^2 s^1: 5s s^0: -9 ?

OpenStudy (anonymous):

yep... so we string them all together with "+" to get...

OpenStudy (anonymous):

descending or ascending order?

OpenStudy (anonymous):

s^5 + -3s^4 + 6s^3 + 12s^2 + 5s + -9 but you can just write a "-" whenever you see a "+ -" descending exponents order...

OpenStudy (anonymous):

can you simplify that then without the "+ -" ?

OpenStudy (anonymous):

S^5 + -3s^4 + 6s^3 +12s^2 + 5s -9

OpenStudy (anonymous):

s^5 - 3s^4 + 6s^3 + 12s^2 + 5s - 9 that's it.. you just forgot the first one... :)

OpenStudy (anonymous):

lol thank you so much!

OpenStudy (anonymous):

yw... be a good daddy.. i'm sure u are... :)

OpenStudy (anonymous):

The whole reason I am spending my weekends doing homework and not drinking beer lol

OpenStudy (anonymous):

nice one...:)